# When Parts Are Greater Than Sums: Individual LLM Components Can Outperform Full Models

### v2

@@ A.4 Tests of Significance
We run one-tailed paired t-tests to test whether CompRW outperforms Calib+ significantly. In Table 4, we have the results of 15 prompts for each task and 8 tasks in total. For each model, we aggregate the 120 accuracy scores of CompRW and Calib+, respectively, and then calculate the p-values. Table 5 shows that p-values $<0.05$ in 6/6 setups, suggesting that CompRW performs significantly better than Calib+.
Llama-2-7B Llama-2-13B Mistral-Instruct-7B
$K=12$ 0.0010 0.0002 0.0470
$K=24$ 0.0003 0.0001 0.0027
Table 5: The p-values $<0.05$ in all 6 setups (3 LLMs, with $K=\{12,24\}$ labeled examples), showing that CompRW performs significantly better than Calib+.
@@ A.5 Do Good-Performing Components Exist in Randomly Initialized Models?
We run 4-shot ICL with 15 prompts and report the average accuracy and standard deviation. For CompRW, we use the same 4 demonstrations and 20 more examples for reweighting. Table 6 shows that the best-performing component in a randomly initialized Llama-2-7B still performs poorly on SST2 and ARC-Easy. While CompRW has substantial improvement over Full on the pretrained model, it has no effect on the randomly initialized model. We conclude that good-performing components do not exist in randomly initialized LLMs and our CompRW method relies on the pretrained component activations to perform well.
@@ A.6 More Results on Transferability
A.7 Training Details and Hyperparameters
@@
Figure 7: Each dot represents a component (attention head: blue; MLP: orange) under 4-shot ICL on Mistral-Instruct-7B. The x-axis shows how often a component predicts label 1 on the test set.
A.8 Implementation Details
We use the model checkpoints on Hugging Face, Llama-2-7b-hf [^4] https://huggingface.co/meta-llama/Llama-2-7b-hf, Llama-2-13b-hf [^5] https://huggingface.co/meta-llama/Llama-2-13b-hf, and Mistral-7B-Instruct-v0.1666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1. We acknowledge that Mistral-7B-Instruct-v0.3 may have better ICL performance than Mistral-7B-Instruct-v0.1; however, it is not supported by TransformerLens [^7] https://github.com/TransformerLensOrg/TransformerLens (the package we use for model decomposition) to date.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC-Easy
Llama-2-7B $37.8$ $18.9$ $43.4$ $44.2$ $35.4$ $28.2$ $13.4$ $11.5$
Llama-2-13B $32.6$ $18.7$ $37.2$ $39.3$ $32.1$ $26.0$ $12.4$ $13.1$
Mistral-Instruct-7B $31.9$ $14.0$ $32.3$ $32.4$ $27.6$ $20.9$ $14.5$ $8.4$
Dataset Task # Classes
SST-2 Socher et al. (2013) Sentiment Analysis 2
Yelp-polarity Zhang et al. (2015) Sentiment Analysis 2
BoolQ Clark et al. (2019) Yes/No QA 2
BoolQ Contrast Set Gardner et al. (2020) Yes/No QA 2
QQP Wang et al. (2018) Paraphrase Identification 2
WiC Pilehvar and Camacho-Collados (2019) Word Sense Disambiguation 2
RTE Wang et al. (2018) Natural Language Inference 2
MNLI Williams et al. (2018) Natural Language Inference 3
MedNLI Romanov and Shivade (2018) NLI in Medical Domain 3
AGNews Zhang et al. (2015) Topic Classification 4
ARC-Easy Clark et al. (2018) Multiple-Choice QA 4
Task069 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Task070 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Table 8: Summary of all the datasets.
@@ A.8 Implementation Details
Table 9: We construct minimally contrastive templates that only differ slightly (colored in red) but yield large differences in 4-shot ICL accuracy on Llama-2-7B. We report the average accuracy and standard deviation across 5 ICL runs under the same template.
@@
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC
Correlation Llama-2-7B
(1) Demo 0.81 0.84 0.60 0.65 0.75 0.65 0.89 0.88
(2) Temp 0.40 0.16 0.03 0.15 0.19 0.09 0.68 0.44
IoU
(1) Demo 0.36 0.74 0.27 0.21 0.53 0.24 0.63 0.70
(2) Temp 0.12 0.01 0.01 0.03 0.05 0.01 0.20 0.20
Correlation Llama-2-13B
(1) Demo 0.83 0.84 0.63 0.67 0.78 0.73 0.91 0.91
(2) Temp 0.57 0.30 0.09 0.19 0.28 0.16 0.76 0.55
IoU
(1) Demo 0.26 0.71 0.31 0.18 0.46 0.39 0.55 0.65
(2) Temp 0.21 0.11 0.07 0.01 0.21 0.07 0.25 0.30
Correlation Mistral-Instruct-7B
(1) Demo 0.88 0.91 0.72 0.75 0.87 0.82 0.92 0.97
(2) Temp 0.58 0.44 0.19 0.26 0.40 0.30 0.77 0.60
IoU
(1) Demo 0.39 0.59 0.27 0.29 0.50 0.45 0.68 0.80
(2) Temp 0.10 0.17 0.06 0.05 0.17 0.09 0.29 0.22
Table 10: Full results of the average correlation and IoU between (1) two random sets of demonstrations and (2) two random templates. ARC is short for the ARC-Easy task.

### v3

@@ A.2 Label-Biased Components
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC-Easy
Llama-2-7B $37.8$ $18.9$ $43.4$ $44.2$ $35.4$ $28.2$ $13.4$ $11.5$
Llama-2-13B $32.6$ $18.7$ $37.2$ $39.3$ $32.1$ $26.0$ $12.4$ $13.1$
Mistral-Instruct-7B $31.9$ $14.0$ $32.3$ $32.4$ $27.6$ $20.9$ $14.5$ $8.4$
Llama-3-8B $38.3$ $21.4$ $42.4$ $40.0$ $36.9$ $28.1$ $15.8$ $15.0$
@@ A.4 Tests of Significance
We run one-tailed paired t-tests to test whether CompRW outperforms Calib+ significantly. In Table 4, we have the results of 15 prompts for each task and 8 tasks in total. For each model, we aggregate the 120 accuracy scores of CompRW and Calib+, respectively, and then calculate the p-values. Table 6 shows that p-values $<0.05$ in 8/8 setups, suggesting that CompRW performs significantly better than Calib+.
Llama2-7B Llama2-13B Mistral-Ins-7B Llama3-8B
$K=12$ 0.0010 0.0002 0.0470 0.0198
$K=24$ 0.0003 0.0001 0.0027 0.0245
Table 6: The p-values $<0.05$ in all 8 setups (4 LLMs, with $K=\{12,24\}$ labeled examples), showing that CompRW performs significantly better than Calib+.
@@ A.5 Do Good-Performing Components Exist in Randomly Initialized Models?
We run 4-shot ICL with 15 prompts and report the average accuracy and standard deviation. For CompRW, we use the same 4 demonstrations and 20 more examples for reweighting. Table 7 shows that the best-performing component (Oracle-T1) in a randomly initialized Llama-2-7B still performs poorly on SST2 and ARC-Easy. While CompRW has substantial improvement over Full on the pretrained model, it has no effect on the randomly initialized model. We conclude that good-performing components do not exist in a randomly initialized LLM and our CompRW method relies on the pretrained component activations to perform well.
@@ A.6 More Results on Transferability
A.7 Pruning Good and Bad Components
Our method studies a component using its cached direct contribution to the output, whereas Michel et al. (2019) (pruning) zeroes out the activations of a component in the forward pass and thus indirectly changes the activations of other components in the upper layers. They consider a component important if pruning it causes large drops in task performance. In this section, we investigate the intersection between our method and pruning.
[…]
SST2 RTE MNLI AGNews
Full Model $\textbf{75.8}_{\hskip 1.42271pt18.1}$ $\textbf{68.9}_{\hskip 1.42271pt3.2}$ $34.4_{\hskip 1.42271pt1.7}$ $70.0_{\hskip 1.42271pt19.9}$
Prune-T50 ${\color[rgb]{1,0,0}53.4}_{\phantom{1}\hskip 1.42271pt8.8}$ ${\color[rgb]{1,0,0}57.8}_{\hskip 1.42271pt6.4}$ $34.3_{\hskip 1.42271pt3.2}$ ${\color[rgb]{1,0,0}26.8}_{\phantom{1}\hskip 1.42271pt2.5}$
Prune-B50 $72.3_{\hskip 1.42271pt14.8}$ $65.4_{\hskip 1.42271pt5.8}$ $\textbf{35.7}_{\hskip 1.42271pt4.0}$ $\textbf{72.6}_{\hskip 1.42271pt15.7}$
Table 8: Comparing the accuracies of the full Llama-2-7B model and pruning the top/bottom 50 components. We run 15 prompts for each task and report the average accuracy and standard deviation. We color the numbers red when there is a large drop in accuracy.
A.8 Training Details and Hyperparameters
@@
A.9 Models
We use the model checkpoints on Hugging Face, meta-llama/Llama-2-7b-hf, Llama-2-13b-hf, mistralai/Mistral-7B-Instruct-v0.1, and meta-llama/Meta-Llama-3-8B.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC
Correlation Llama-2-7B
(1) Demo 0.81 0.84 0.60 0.65 0.75 0.65 0.89 0.88
(2) Temp 0.40 0.16 0.03 0.15 0.19 0.09 0.68 0.44
IoU
(1) Demo 0.36 0.74 0.27 0.21 0.53 0.24 0.63 0.70
(2) Temp 0.12 0.01 0.01 0.03 0.05 0.01 0.20 0.20
Correlation Llama-2-13B
(1) Demo 0.83 0.84 0.63 0.67 0.78 0.73 0.91 0.91
(2) Temp 0.57 0.30 0.09 0.19 0.28 0.16 0.76 0.55
IoU
(1) Demo 0.26 0.71 0.31 0.18 0.46 0.39 0.55 0.65
(2) Temp 0.21 0.11 0.07 0.01 0.21 0.07 0.25 0.30
Correlation Mistral-Instruct-7B
(1) Demo 0.88 0.91 0.72 0.75 0.87 0.82 0.92 0.97
(2) Temp 0.58 0.44 0.19 0.26 0.40 0.30 0.77 0.60
IoU
(1) Demo 0.39 0.59 0.27 0.29 0.50 0.45 0.68 0.80
(2) Temp 0.10 0.17 0.06 0.05 0.17 0.09 0.29 0.22
Correlation Llama-3-8B
(1) Demo 0.85 0.88 0.70 0.73 0.80 0.81 0.89 0.95
(2) Temp 0.55 0.39 0.26 0.25 0.31 0.23 0.67 0.52
IoU
(1) Demo 0.42 0.56 0.28 0.25 0.46 0.52 0.65 0.68
(2) Temp 0.15 0.12 0.09 0.07 0.08 0.05 0.34 0.27
Table 9: Full results of the average correlation and IoU between (1) two random sets of demonstrations and (2) two randomly sampled templates.
@@ A.9 Models
Table 10: We construct minimally contrastive templates that only differ slightly (colored in red) but yield large differences in 4-shot ICL accuracy on Llama-2-7B. We report the average accuracy and standard deviation across 5 ICL runs with different demonstrations under the same template.
Dataset Task # Classes
SST-2 Socher et al. (2013) Sentiment Analysis 2
Yelp-polarity Zhang et al. (2015) Sentiment Analysis 2
BoolQ Clark et al. (2019) Yes/No QA 2
BoolQ Contrast Set Gardner et al. (2020) Yes/No QA 2
QQP Wang et al. (2018) Paraphrase Identification 2
WiC Pilehvar and Camacho-Collados (2019) Word Sense Disambiguation 2
RTE Wang et al. (2018) Natural Language Inference 2
MNLI Williams et al. (2018) Natural Language Inference 3
MedNLI Romanov and Shivade (2018) NLI in Medical Domain 3
AGNews Zhang et al. (2015) Topic Classification 4
ARC-Easy Clark et al. (2018) Multiple-Choice QA 4
Task069 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Task070 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Table 11: Summary of all the datasets.
@@
Figure 8: Each dot represents a component (attention head: blue; MLP: orange) under 4-shot ICL on Mistral-Instruct-7B. The x-axis shows how often a component predicts label 1 across the test data of a binary task.